Two point charges are located on the -axis of a coordinate system: q1 = 1.0 nC is at x = +2.0 cm, and q2 is at x = +4.0 cm. What is the total electric force exerted by q1 and q2 on a charge q3 = 5.0 nC at x = 0?


1
Expert's answer
2022-02-16T08:35:51-0500

Answer

Net Force on q3 is in same direction due to q1 and q2 ,so these added up

"F=kq_3(\frac{q_1}{(0.02)^2}+\frac{q_2}{(0.04)^2}) \\=9*10^9*5*10^{-9}(\frac{1*10^{-9}}{(0.02)^2}+\frac{1*10^{-9}}{(0.04)^2}) \\=45(2500+625)*10^{-9}\\=0.14*10^{-3}N"